# Are All Marine Species Created Equal? Performance Disparities in Underwater Object Detection

### v1

@@ Abstract
Underwater object detection is critical for monitoring marine ecosystems but poses unique challenges, including degraded image quality, imbalanced class distribution, and distinct visual characteristics. Not every species is detected equally well, yet underlying causes remain unclear. We address two key research questions: 1) What factors beyond data quantity drive class-specific performance disparities? 2) How can we systematically improve detection of under-performing marine species? We manipulate the DUO dataset to separate the object detection task into localization and classification and investigate the under-performance of the scallop class. Localization analysis using YOLO11 and TIDE finds that foreground-background discrimination is the most problematic stage regardless of data quantity. Classification experiments reveal persistent precision gaps even with balanced data, indicating intrinsic feature-based challenges beyond data scarcity and inter-class dependencies. We recommend imbalanced distributions when prioritizing precision, and balanced distributions when prioritizing recall. Improving under-performing classes should focus on algorithmic advances, especially within localization modules. We publicly release our code and datasets.
@@ 1 Introduction
Oceans are crucial for the regulation and function of natural water and carbon cycles, so much so that all life on Earth relies on healthy marine ecosystems. However these ecosystems are under threat from anthropogenic impacts including pollution, over-fishing and climate change [25]. This increases the urgency and importance of monitoring these environments.
@@
Figure 1: Example underwater image from the DUO dataset [17], showing the four target classes (top row): Holothurian, Starfish, Scallop and Echinus. Left middle: the original class distribution of DUO shown by the lighter colored bars on top, and the DUO dataset with down-sampled, class-balanced distribution in the solid bars below. Left bottom: localization performance (mAP@0.5) on single-class versions of DUO. The lightened top bars indicate the localization performance for each class under the original distribution, and the solid colored bar below demonstrates there is still a clear performance difference between the classes, even when there is an equal number of training instances.
However, object detection in underwater imagery is significantly more challenging than for terrestrial images. Not only is the image quality often degraded by turbidity, blur and low contrast [32], but there also are certain visual characteristics and behavioral habits associated with different marine species that complicate their reliable detection. Image datasets of marine species may exhibit class imbalance, varying sizes of individuals, intra-class variation in appearance, background camouflaging, clustering of individuals, and dense habitats. Naturally, this leads to a discrepancy in the object detection performance for different classes of marine species (Fig. 1).
The Detecting Underwater Objects (DUO) dataset [17] is a combination of six earlier datasets, and contains real seafloor images of scallops, sea urchins (Echinus), sea cucumbers (Holothurian) and starfish [17], with an example image seen in Fig. 1. Prior underwater object detection (UOD) works [33, 31, 6, 32, 10, 34, 15] use this dataset as a benchmark, and demonstrate a clear trend in the relative performance of the different classes (Tab. 1). In these works, the authors attribute the significant under-performance of the scallop class in particular to the scarcity of instances in the training data [17, 33, 10] or the performance difference is not addressed at all.
@@
Building on these observations, this work does not propose a new architecture but instead provides a systematic analysis of the under-performance of specific species in underwater object detection, moving beyond the commonly cited explanation of data scarcity. We investigate two important factors: the dataset characteristics, specifically the class distribution and data quantity; and the influence of inherent visual characteristics of the target objects. While the former can be changed when collecting the data and choosing the samples, the latter needs to be addressed on the algorithmic level. We also decompose the object detection architecture into the core tasks of localization and classification, to examine which stages are more impacted by these different data challenges. Understanding the underlying reasons for the class-wise detection performance is crucial to design improved object detectors for underwater ecosystems.
@@
We make our evaluation and data processing code available to foster further research in this area [^1] Link will be added upon acceptance..
@@ 2.1 Challenges of the Underwater Domain
Transferring generic object detection to marine environments poses unique challenges caused by the characteristics of underwater imagery; these include poor visibility, turbidity, blur, low contrast, light scattering, and color distortion, and complicate feature extraction and foreground-background distinction [9, 25, 32, 8]. These challenges are compounded by the small size of many marine targets, frequent clustering of targets, lack of visual features, occlusion, and unclear object boundaries [8, 10, 15]. Few researchers consider the natural camouflage ability of marine species [8].
Learning the distinct features of marine species is therefore difficult and further complicated by in-class variation in appearance and scale, and differences in the angle, distance and speed of the camera [32, 25, 2]. While abundant underwater images can be obtained, the annotation effort is very time and cost-intensive and requires domain experts to reduce the risk of noisy and incorrect annotations [25, 2]. Finally, class imbalance is a common issue in underwater data, and is caused by differences in the economic value and rarity of certain species [18, 31].
@@ 2.3 Failure Analysis
However, if only the aggregated metrics are used, the localization accuracy (i.e. how well bounding boxes are placed) and potential performance discrepancies across classes remain hidden. The respective influence of False Negative (FN) and False Positive (FP) errors are ignored. Hence different precision-recall curves can yield the exact same AP score, despite fundamentally different error patterns [23]. A detector might have low recall but high precision, making it suitable for risky tasks like autonomous underwater navigation. Conversely, another detector with the same mAP might instead provide high recall but low precision, which is preferred for ecological monitoring of invasive and harmful species, when it is most critical to avoid missing occurrences and enable early warnings.
@@ 3 Experimental Framework
This section details the general baseline for all our experiments, summarizing our approach, introducing the dataset of interest and the performance metrics used.
Approach: Object detection can be broadly considered as 3 key stages: (i) distinguishing between the background of an image and objects of interest; (ii) once detected as object of interest, tightly localizing it with a bounding box; and (iii) once localized as object of interest, classifying it between target classes. We investigate each of these in turn to identify how class-wise performance is affected by inherent visual features versus characteristics of the dataset. Steps (i) and (ii) are part of our localization analysis (Sec. 4), whereas (iii) is examined in a classification-focused study (Sec. 5). In every step we consider the influence of data quantity and distribution versus inherent class-specific properties.
Base Dataset: We perform our experiments on data from the DUO dataset [17]. DUO provides a standardized benchmark and assembles images from the Underwater Robot Professional Contest from 2017 to 2020, as well as the Underwater open-sea farm object Detection Dataset (UDD) [18]. In its original form, DUO provides 6,671 training and 1,111 test images, with 50,156 labeled echinus, 7,887 holothurian, 1,924 scallop and 14,548 starfish detections. We generate several subsets from the original DUO for our localization and classification studies, as detailed in Secs. 4.2 and 5.2.
Evaluation Metrics: We use mAP for the localization experiments and precision and recall for classification.
We use the common object detection metric AP for the localization task, with $\text{AP}=\int_{0}^{1}p(r)\,dr$. When averaged over the number of classes $C$, this becomes $\text{mAP}=\frac{1}{C}\sum_{i=1}^{C}\text{AP}_{i}$.
@@
Precision $p=\frac{TP}{(TP+FP)}$ is the proportion of correct predictions (true positives $TP$) out of all predictions made. Recall, also known as the True Positive Rate (TPR), $\text{TPR}=\frac{TP}{(TP+FN)}$ is the proportion of correctly predicted objects out of all true objects. The False Negative Rate (FNR) is the proportion of missed targets out of the total, $\text{FNR}=\frac{FN}{(TP+FN)}$. We also consider the False Discovery Rate, which is the ratio of false positives to total predictions, $\text{FDR}=\frac{FP}{(FP+TP)}$. Finally, the F1 Score balances precision and recall: $F1=2\cdot\frac{\text{Precision}\cdot\text{Recall}}{\text{Precision}+\text{Recall}}$.
@@ 4 Localization Study
This section describes our localization analysis by explaining the methodology (Sec. 4.1), providing an overview of the datasets (Sec. 4.2), and detailing the implementation (Sec. 4.3). We then provide our results that show how well each class can be distinguished from the background (Sec. 4.4) and how well the object can be located with a bounding box (Sec. 4.5).
@@ 4.1 Method
In our localization experiments, we eliminate the classification stage interference by creating single-class versions of the DUO dataset. We train object detectors on individual classes to analyze class-specific localization challenges independently. The focus lies on assessing foreground-background confusion and examining bounding box quality. We also leverage TIDE [1] to gain detailed insights into the underlying error types.
@@ 4.2 Datasets
We obtain single-class DUO versions using the original DUO images but modify the annotations to omit the labels for all but one species at a time. A random 20% from the training split is selected for the validation split. To obtain further subsets, we select all “empty” images first and then add images containing the target species until reaching the desired instance count. Since images may contain multiple objects of the same class, we include all instances of the target species when an image is selected. The subsets include: (i) balanced sets in which we use the minority class (scallop) count as the target: 1,363 instances for training and 344 for validation, so that each single-class dataset contains an equal number of training objects; (ii) reduced sets with the instance count decreased to 75%, 50% and 25% of the balanced amount in each single-class dataset; and (iii) a special single-class DUO dataset (limited scallop) which consists only of images that actually contain the target of interest, scallop, for comparison purposes.
@@ 4.3 Implementation
Figure 3: Per class localization performance compared across datasets. From the left: mAP@0.5 for the single class DUO versions and balanced sets, which are then gradually reduced. Scallop performance is the same in the original DUO and balanced sets (all classes in the balanced set are downsampled to the scallop instance count), but still lower than all other classes, despite the same number of instances. Further reductions of the balanced set affect classes differently: starfish and echinus are impacted less by the decrease in samples, whereas holothurian and scallop classes exhibit significant reduction in the mAP.
For the localization experiments, we used Ultralytics YOLO11222https://github.com/ultralytics/ultralytics as object detector. The model was loaded with the default parameters, that is, 30 epochs, image resized to 640x640 pixels and data augmentations during training, including hue, saturation, brightness, translation, image scaling, horizontal flip and mosaic augmentation. Batch size was set to 8. The best model is saved based on the validation performance and evaluated on the test set with confidence threshold 0.25 and IoU set to 0.7. All localization experiments were run on an NVIDIA H100 GPU.
@@ 4.4 Results: Foreground-Background Separation
Our localization experiments reveal that class performance disparities persist even when the class distribution in the DUO dataset is balanced. As the left side of Fig. 3 illustrates, the single-class DUO versions show the drastic performance gaps that have motivated our work – mAP@0.5 scores for echinus and starfish greater than 90%, followed by holothurian at 82.5%, but substantially lower scallop performance at 44% – yet, even when all datasets are balanced to contain the same number of training labels, these performance gaps remain evident. The difference between scallop and holothurian for mAP@0.5 (the smallest gap) remains 27%, clearly demonstrating that the model struggles with scallop localization regardless of the quantity of training samples. This is the most major performance difference of all three detector tasks, and hints at challenges due to the inherent visual features.
Fig. 3 (right) shows that when the data quantity is gradually decreased in our balanced setup, the holothurian and scallop classes are more affected by the number of sample points in the training dataset and sensitive to data reductions. While the downwards trend generally highlights the benefits of larger training datasets, the varying rates of performance change again hint at inherent data characteristic differences.
single-class DUO balanced single-class sets
@@
Holothurian 76.74 21.89 23.26 56.81 19.24 43.19
Echinus 87.97 14.16 12.03 51.80 4.04 48.20
Starfish 88.02 16.95 11.98 68.32 8.42 31.68
Scallopall images 23.96 35.8 76.04 same as DUO
Scalloplimited 74.19 75.57 25.81
@@
Table 2: True Positive Rate (TPR) or Recall, False Discovery Rate (FDR), and False Negative Rate (FNR) for each class under two training setups (refer to Sec. 3 for metric definitions). Cell colors indicate the relative contribution. Most of the scallop instances are undetected (76% FNR). Despite the increases in FN for Holothurian, Echinus and Starfish on the balanced set, their recall remains above the scallop value (24%). Training without background-only images (Scalloplimited row) increases the recall by 50.2%, but at the expense of an increase of 39.8% in the FDR.
(a) Holothurian DUO
(b) Echinus DUO
(c) Starfish DUO
(d) Scallop DUO = Balanced
(e) Holothurian Balanced
(f) Echinus Balanced
(g) Starfish Balanced
(h) Scallop Limited
@@
Figure 4: Distribution of TIDE error types across all classes for both original and balanced DUO datasets. The imbalanced versions exhibit higher rates of FPs, and specifically background error, for Starfish, echinus and holothurian classes. In contrast, balanced datasets make the error profiles for these classes more similar to the scallop class, i.e. there are more missed ground truth detections. Only training on object-containing images (as seen in the (h) Scallop Limited sub-plot) reduces misses but massively increases background error.
@@
Tab. 2 breaks down the true positive (recall), false discovery and false negative rates under imbalanced and balanced conditions. The results reveal that scallops generally suffer from extremely poor recall, with only 24% of true scallops being detected, compared to other classes with recall above 75% for single-class DUO. The balanced single-class sets resulted in decreased recall by 19.9%, 36.2% and 19.7% for holothurian, echinus and starfish, respectively. However, despite the equal number of training instances across the classes, scallops still show significantly lower recall at 24% than the other classes (51.8-68.3%), and are affected significantly more by FNs (76% for scallop compared with 31.7-48.2%), i.e. they are most often mistaken as part of the background. In the single-class DUO sets, the FDR is also highest for the scallop class at 35.8%, as compared to 14-21% for the other classes. When the sets are balanced, the FDR reduces by 2.7% to 19.2% for holothurian, by 10.1% to 4.0% for echinus, and by 8.5% to 8.4% for starfish, increasing the gap to the high scallop FDR to a maximum difference of 16.6%.
@@
Qualitative inspection confirms the high background confusion, with scallops being the hardest to spot due to their ability to camouflage. In Fig. 2, the missed example is a small scallop blending into the background, resembling a faint light spot that could easily be mistaken for scattered light or a rock. The background error shows a rock falsely detected as a scallop, highlighting their visual similarity. In contrast, other classes are more distinct –echinus are dark and spiny, starfish have clear five-arm symmetry, and holothurians are elongated with surface spikes. Thus, scallops are arguably the most difficult class to distinguish from background clutter.
@@
Overall, background confusion is the most prominent TIDE error type that occurs across all classes, as visible in Fig. 4. With the original DUO instances per class, starfish, echinus and holothurian still result in a considerable amount of background errors (causing FPs), the balanced versions result in high rates of missed GT errors. Of all the classes, scallops have the highest error rate relative to the number of test objects.
@@
An important consideration when creating our datasets was whether to include background-only images. The comparisons between full scallop datasets and limited scallop-only datasets in Tab. 2 and Fig. 4 reveal that including diverse background images (even those without scallops) improves model generalization and heavily reduces false positives, though recall is reduced compared to training only on scallop-containing images.

### v2

@@ Abstract
Underwater object detection is critical for monitoring marine ecosystems but poses unique challenges, including degraded image quality, imbalanced class distribution, and distinct visual characteristics. Not every species is detected equally well, yet underlying causes remain unclear. We address two key research questions: 1) What factors beyond data quantity drive class-specific performance disparities? 2) How can we systematically improve detection of under-performing marine species? We manipulate the DUO and RUOD datasets to separate the object detection task into localization and classification and investigate the under-performance of the scallop class. Localization analysis using YOLO11 and TIDE finds that foreground-background discrimination is the most problematic stage regardless of data quantity. Classification experiments reveal persistent precision gaps even with balanced data, indicating intrinsic feature-based challenges beyond data scarcity and inter-class dependencies. We recommend imbalanced distributions when prioritizing precision, and balanced distributions when prioritizing recall. Improving under-performing classes should focus on algorithmic advances, especially within localization modules. We publicly release our code and datasets.
@@ 1 Introduction
Oceans are crucial for the regulation and function of natural water and carbon cycles, making healthy marine ecosystems essential for life on Earth. However these ecosystems are under threat from anthropogenic impacts including pollution, over-fishing and climate change [28]. This increases the urgency and importance of monitoring these environments.
@@
Figure 1: Example underwater image from the DUO dataset [19], showing the four target classes (top row): Holothurian, Starfish, Scallop and Echinus. Left middle: the original class distribution of DUO shown by the lighter colored bars on top, and the DUO dataset with down-sampled, class-balanced distribution in the solid bars below. Left bottom: localization performance (mAP@0.5) on single-class versions of DUO. The lightened top bars indicate the localization performance for each class under the original distribution, and the solid colored bars below demonstrates there is still a clear performance difference between the classes, even when there is an equal number of training instances.
However, object detection in underwater imagery is significantly more challenging than for terrestrial images. Not only is the image quality often degraded by turbidity, blur and low contrast [36], but there are certain visual characteristics and behavioral habits associated with different marine species that complicate their reliable detection. Image datasets of marine species may exhibit class imbalance, varying sizes of individuals, intra-class variation in appearance, background camouflaging, clustering of individuals, and dense habitats. Naturally, this leads to a discrepancy in the object detection performance for different classes of marine species (Fig. 1).
The Detecting Underwater Objects (DUO) dataset [19] is a combination of six earlier datasets, and contains real seafloor images of scallops, sea urchins (Echinus), sea cucumbers (Holothurian) and starfish [19], with an example image seen in Fig. 1. Prior underwater object detection (UOD) works [37, 34, 7, 36, 11, 38, 17] use this dataset as a benchmark, and demonstrate a clear trend in the relative performance of the different classes (Tab. 1). In these works, the authors attribute the significant under-performance of the scallop class in particular to the scarcity of instances in the training data [19, 37, 11] or the performance difference is not addressed at all. A similar class-specific performance pattern can be observed in the Rethinking general Underwater Object Detection (RUOD) dataset [12], which was carefully collected from various photo-sharing websites to provide data with challenging environmental conditions and a diverse range of ten marine species, including the common categories also contained in DUO. We focus on these four categories and examine some prior studies on the RUOD dataset that report per-class metrics [3, 21, 35], finding that these works also show a performance gap between scallops and the other classes, although this gap is not as pronounced as in DUO. As for DUO, these works do not provide sufficient analysis nor justification for the scallop under-performance.
@@
Building on these observations, this work presents the first systematic analysis into the under-performance of specific species in underwater object detection. We do not propose a novel architecture in this work, and instead identify common failure modes and species-specific factors, moving beyond the often-cited explanation of data scarcity, and offer key recommendations to guide future research in this area. We investigate two important factors: the dataset characteristics, specifically the class distribution and data quantity; and the influence of inherent visual characteristics of the target objects. While the former can be changed when collecting the data and choosing the samples, the latter needs to be addressed on the algorithmic level. We also decompose the object detection architecture into the core tasks of localization and classification, to examine which stages are more impacted by these different data challenges. Understanding the underlying reasons for the class-wise detection performance is crucial to design improved object detectors for underwater ecosystems.
@@
We will make our evaluation and data processing code available upon acceptance to foster further research in this area.
@@ 2.1 Challenges of the Underwater Domain
Transferring generic object detection to marine environments poses unique challenges caused by the characteristics of underwater imagery; these include poor visibility, turbidity, blur, low contrast, light scattering, and color distortion, and complicate feature extraction and foreground-background distinction [10, 28, 36, 9]. These challenges are compounded by the small size of many marine targets, frequent clustering of targets, lack of visual features, occlusion, and unclear object boundaries [9, 11, 17]. Few researchers consider the ability of marine species to camouflage [9, 12].
Learning the distinct features of marine species is therefore complicated by in-class variation in appearance and scale, and differences in the angle, distance and speed of the camera [36, 28, 2]. While abundant underwater images can be obtained, the annotation effort is very time and cost-intensive and must be performed by domain experts to reduce the risk of noisy and incorrect annotations [28, 2]. Finally, class imbalance is a common issue in underwater data, and is caused by differences in the economic value and rarity of certain species [20, 34].
@@ 2.3 Failure Analysis
However, if only the aggregated metrics are used, the localization accuracy and potential performance discrepancies across classes remain hidden. The respective influence of False Negative (FN) and False Positive (FP) errors are ignored. Hence different precision-recall curves can yield the exact same AP score, despite fundamentally different error patterns [26]. A detector might have low recall but high precision, making it suitable for risky tasks like autonomous underwater navigation. Conversely, another detector with the same mAP might instead provide high recall but low precision, which is preferred for ecological monitoring of invasive and harmful species, when it is most critical to avoid missing occurrences and to enable early warnings.
@@ 3 Experimental Framework
This section details the general baseline for all our experiments bysummarizing our approach, introducing the dataset of interest and the performance metrics used.
Approach: Object detection can be broadly considered as three key stages: (i) distinguishing between the background of an image and objects of interest; (ii) detected objects of interest are tightly localized with a bounding box; and (iii) localized objects of interest are classified. We investigate each of these in turn to identify how class-wise performance is affected by inherent visual features versus characteristics of the dataset. Steps (i) and (ii) are part of our localization analysis (Sec. 4), whereas (iii) is examined in a classification-focused study (Sec. 5). In every step we consider the influence of data quantity and distribution versus inherent class-specific properties.
Base Dataset: We perform our experiments with a primary focus on the DUO dataset [19]. DUO provides a standardized benchmark and assembles images from the Underwater Robot Professional Contest from 2017 to 2020, as well as the Underwater open-sea farm object Detection Dataset (UDD) [20]. In its original form, DUO provides 6,671 training and 1,111 test images, with 50,156 labeled echinus (67%), 7,887 holothurian (11%), 1,924 scallop (3%) and 14,548 starfish objects (19%). We generate several subsets from the original DUO for our localization and classification studies, as detailed in Secs. 4.2 and 5.2. To confirm the generalizability of the results, we also apply our experimental framework to a modified version of the RUOD dataset [12], called RUOD-4C, which only includes the images that contain echinus, holothurian, scallop and/or starfish annotations, i.e. 11,292, 7,626, 7,411 and 8,118 objects. The RUOD-4C class distribution is relatively even (33%, 22%, 22%, 23% respectively). While the original RUOD consists of 9,800 training and 4,200 testing images, these numbers are reduced to 3,013 and 1,255 images in our RUOD-4C.
Evaluation Metrics: We use the common object detection metric mAP for the localization experiments, and precision and recall for classification. The mAP is calculated by taking the average precision, $\text{AP}=\int_{0}^{1}p(r)\,dr$, and averaging over the classes $C$: $\text{mAP}=\frac{1}{C}\sum_{i=1}^{C}\text{AP}_{i}$.
@@
To evaluate classification performance, we look at the proportion of correct predictions (true positives, $TP$) out of all predictions made, which is given as precision $p=\frac{TP}{(TP+FP)}$, and the proportion of correctly predicted objects out of all true objects, called recall or True Positive Rate (TPR), $\text{TPR}=\frac{TP}{(TP+FN)}$. We also consider the False Negative Rate (FNR), as the proportion of missed targets out of the total, $\text{FNR}=\frac{FN}{(TP+FN)}$, and the False Discovery Rate, which is the ratio of false positives to total predictions, $\text{FDR}=\frac{FP}{(FP+TP)}$. Finally, the F1 Score balances precision and recall: $F1=2\cdot\frac{\text{Precision}\cdot\text{Recall}}{\text{Precision}+\text{Recall}}$.
@@ 4 Localization Study
This section describes our localization analysis by providing an overview of the methodology (Sec. 4.1), datasets (Sec. 4.2) and implementation details (Sec. 4.3). We then discuss our results evaluating how well each class is distinguished from the background (Sec. 4.4) and our results evaluating bounding box localization (Sec. 4.5).
@@ 4.1 Method
In our localization experiments, we eliminate the classification stage interference by creating single-class versions of the datasets and train object detectors on them to analyze class-specific localization challenges independently. The focus lies on assessing foreground-background confusion and examining bounding box quality. We also leverage TIDE [1] to gain detailed insights into the underlying error types.
@@ 4.2 Datasets
We obtain single-class DUO/RUOD-4C versions using the original images but modifying the annotations to omit the labels for all but one species at a time. A random 20% from the training split is selected for the validation split. To obtain further subsets, we select all “empty” images first and then add images containing the target species until reaching the desired instance count. Since images may contain multiple objects of the same class, we include all instances of the target species when an image is selected. The subsets include: (i) balanced sets in which we use the minority class (scallop) count as the target, so that each single-class dataset contains an equal number of training objects; (ii) reduced sets with the instance count decreased to 75%, 50% and 25% of the balanced amount in each single-class dataset; and (iii) a special single-class DUO dataset (limited scallop) which does not include any “empty” images but consists only of images that actually contain the target of interest, scallop, for comparison purposes. This is specific to DUO as the class-imbalance is more pronounced and hence, the difference between the sets is greater.
@@ 4.3 Implementation
(a) DUO
(b) RUOD-4C
Figure 3: Per class localization performance compared across datasets. Bar charts indicate mAP@0.5 for the originally imbalanced and balanced data, line charts represent gradually reduced sets. For both DUO (a) and RUOD-4C (b), scallop performance is the same in the original and balanced sets (all classes in the balanced set are downsampled to the scallop instance count), but still lower than all other classes, despite the same number of instances. Further reductions of the balanced set affect classes differently: starfish and echinus are impacted less by the decrease, whereas holothurian and scallop exhibit significant reduction in mAP.
For the localization experiments, we used Ultralytics YOLO11n [^1] https://github.com/ultralytics/ultralytics as the object detector, pretrained on COCO. The best model is saved based on the validation performance and evaluated on the test set with confidence threshold 0.25 and IoU set to 0.7 (refer to Supp. Mat. for additional implementation details). To verify that our findings are not specific to YOLO11n, we repeat the localization study with an SSD detector and obtained similar results, as detailed in the Supp. Mat.
@@ 4.4 Results: Foreground-Background Separation
Our localization experiments reveal that class performance disparities persist even when the class distribution is balanced. As the left side of Fig. 3(a) illustrates, the single-class DUO versions show the drastic performance gaps that have motivated our work – mAP@0.5 scores for echinus and starfish greater than 90%, followed by holothurian at 82.5%, but substantially lower scallop performance at 44% – yet, even when all datasets are balanced to contain the same number of training labels, these performance gaps remain evident. The difference between scallop and holothurian for mAP@0.5 (the smallest gap) remains 27%, clearly demonstrating that the model struggles with scallop localization regardless of the quantity of training samples. This is the most major performance difference of all three detector tasks, and hints at challenges due to the inherent visual features.
The right half of Fig. 3(a) shows that when data quantity is gradually decreased in our balanced setup, the holothurian and scallop classes are more affected by the number of sample points in the training data and sensitive to reductions. While the downwards trend generally highlights the benefits of larger training datasets, the varying rates of performance change again hint at inherent data characteristic differences.
Fig. 3(b) confirms the same results on RUOD-4C. Although the scallop under-performance is less pronounced with a 7% gap between balanced holothurian and scallop mAPs, all insights gained from DUO remain valid.
Original Distribution (Imbalanced) Balanced Sets
@@
Ho 76.7/78.7 21.9/27.8 23.3/21.3 56.8/76.4 19.2/25.1 43.2/23.6
Ec 88.0/89.5 14.2/30.1 12.0/10.5 51.8/82.6 4.0/30.8 48.2/17.4
St 88.0/87.8 17.0/23.1 12.0/12.2 68.3/85.6 8.4/19.1 31.7/14.4
Sc 24.0/67.1 35.8/30.7 76.0/32.9 24.0/67.1 35.8/30.7 76.0/32.9
Sclim 74.2/ —– 75.6/ —– 25.8/ —–
@@
Table 2: DUO / RUOD-4C True Positive Rate (TPR), False Discovery Rate (FDR), and False Negative Rate (FNR) per class under two training setups (refer to Sec. 3 for metric definitions). Most scallop instances are undetected (76% FNR). Despite increases in FN for Holothurian, Echinus and Starfish on the balanced set, their recall remains above the scallop value (24%). Training without background-only images (Scalloplimited row) increases the recall by 50.2%, but at the expense of an increase of 39.8% in the FDR.
(a) Holothurian
(b) Echinus
(c) Starfish
(d) Scallop
@@
Figure 4: Distribution of TIDE error types across all classes for the original (red) and balanced (green) DUO datasets. The imbalanced versions exhibit higher rates of FPs, and specifically background error, for Starfish, echinus and holothurian classes. In contrast, balanced datasets make the error profiles for these classes more similar to the scallop class, i.e. there are more missed ground truth detections. Only training on object-containing images, as seen in the Scallop Limited bars (blue), reduces misses but massively increases background error.
@@
Tab. 2 breaks down the true positive (recall), false discovery and false negative rates under imbalanced and balanced conditions. The results reveal that scallops generally suffer from the worst recall and accordingly are most affected by FNs. In DUO, recall is extremely poor with only 24% of true scallops being detected, compared to other classes with recall above 75% for imbalanced data. The balanced sets resulted in decreased recall by 19.9%, 36.2% and 19.7% for holothurian, echinus and starfish, respectively, but even that performance decline is not enough to close the gap. Despite the equal number of training instances, the scallop FN rate (76%) is much higher compared to the other classes, where only 31.7-48.2% are being missed i.e. scallops are most often mistaken as part of the background. In the imbalanced DUO sets, the FDR is also highest for the scallop class at 35.8%, as compared to 14-21% for the other classes. When the sets are balanced, the FDR reduces by 2.7% to 19.2% for holothurian, by 10.1% to 4.0% for echinus, and by 8.5% to 8.4% for starfish, increasing the gap to the high scallop FDR to a maximum difference of 16.6%.
In RUOD-4C, the performance disparities are not as strong but are still evident. The FDRs are generally higher in RUOD-4C than in DUO, with scallops and holothurians being prone to false positive detections most. The scallop class notably shows the lowest recall and highest FNR in both class distributions.
@@
Qualitative inspection confirms the high background confusion, with scallops being the hardest to spot due to their ability to camouflage. In Fig. 2, the missed example is a small scallop blending into the background, resembling a faint spot, easily mistaken for scattered light or a rock. The background error shows a rock falsely detected as a scallop, highlighting their visual similarity. In contrast, other classes are more distinct – echinus are dark and spiny, starfish have clear five-arm symmetry, and holothurians are elongated with surface spikes. Thus, scallops are arguably the most difficult class to distinguish from background clutter.
@@
Overall, background confusion i.e. misinterpreting background as an object or an object as background, is the biggest issue. This is evident Fig. 4, highlighing background errors (Bkg) that cause FPs and missed GT errors (Miss) causing FNs as the most prominent error types that occur across all classes. With the original DUO instance count per class, starfish, echinus and holothurian still result in a considerable amount of Bkg’s, while the balanced versions yield high rates of missed GT errors. Of all the classes, scallops have the highest error rate relative to the number of test objects.
@@
An important consideration when creating our datasets was whether to include background-only images. The comparisons between full scallop datasets and limited scallop-only datasets of DUO in Tab. 2 and Fig. 4 reveal that including diverse background images (even those without scallops) improves model generalization and heavily reduces false positives, though recall is reduced compared to training only on scallop-containing images.